In 2018, Variman, Incorporated had Gross Accounts receivable of $6,456,100 and management estimated the Allowance for Doubtful accounts to be $320,600. Compute the ratio of Allowance of Doubtful accounts to Net Accounts receivable for Variman, Inc. for 2018 (that is, compute Allowance/Net Accounts receivable – record your answer as a percent rounded to one decimal place; for example, record 0.234567 as 23.5%).

In 2018, the Whaddock Company purchased 745,000 units from its supplier at a cost of $160.00 per unit. Whaddock sold 625,000 units of its product in 2018 at a price of $220.00 per unit. Whaddock began 2018 with 125,000 units in inventory (inventory is carried at a cost of $160.00 per unit). Using this information, compute Whaddock’s gross profit for 2018.

Referring back to the previous question, compute Whaddock’s 2018 ending inventory balance in dollars (valued at cost per unit of $160.00).
Answer #1

1

Net accounts receivable = Gross accounts receivable – allowance for doubtful accounts.

=6,456,100 -320,600

=$6135500

ratio of Allowance of Doubtful accounts to Net Accounts receivable=320,600/6135500=5.23%

2

Whaddock’s gross profit for 2018.

Sales-Cost of good sold=gross profit

gross profit=(625000)*($220-$160)=$37,500.000

ending inventory balance in dollars=Beginning Inventory+Purchase Less Units Sold

=125000 units +745000 units -625000=245000 units